OpenStudy (anonymous):

(x-2)^2 + (y-2)^2=4 is a circle. Write the formula for the upper half of the circle. i.e. y=f(x) and what is the domain

OpenStudy (anonymous):

Foil everything out

OpenStudy (anonymous):

x^2-4x+4+y^2-4y-4=4

OpenStudy (anonymous):

x^2-4x+y^2-4y=4

OpenStudy (anonymous):

Solve for y y=sqrt(4-(x-2)^2)+2 You get a plus/minus in the square root but you want the top half so make it positive.

OpenStudy (anonymous):

The the domain is where 4-(x-2)^2>=0 4>=(x-2)^2 +/-2>=x-2 x>=2+/-2 2>=x>=0

OpenStudy (anonymous):

hm i always thought the domain was a coordinate